OpenStudy (anonymous):

Peyton has a credit card with an annual rate of 24.7% compounded monthly. She used the credit card to purchase cleaning supplies in the amount of $189.56. She can pay up to $72 on the credit card each month. How much total interest will she pay?

OpenStudy (anonymous):

rate of interest = 24.7/12 % compounded monthly 1st month : Peyton pays $72. Remaining amount= 189.56-72 = $ 117.56 interest charged = 117.56*(24.7/12)*1/100 =$ 2.42 Now she owes 117.56 + 2.42 = $ 119.98 2nd month: Peyton pays $72 Remaining amount = 119.98 -72 = $ 47.98 interest charged = 47.98*(24.7/12)*1/100 =$ 0.99 Now she owes $ 48.97 3rd Month: She pays $48.97 and now she owes nothing. TOTAL INTEREST PAID = $ (72*2 + 48.97) - 189.56 = $ 3.41

OpenStudy (anonymous):

Wrong